1. (10 points) A company reports the following stockholders equity: Common stock, $2 par, 5,000,000 shares authorized Paid-in capital in excess of par, Common stock Total paid-in capital Retained earnings Total stockholders equity 1300,000 $4,300,000 し400.000 $5.700,000 Compute the (1) number of common shares outstanding and (2) book value per common share. 2. (10 points) On August 1, a companys board of directors declared a 10% stock dividend to be distributed on September 1 to the stockholders of record on August 20. The company had 1,000,000 shares of $2.50 par value common stock outstanding with a market value of $23 per share. Prepare the journal entries required on August 1, August 20, and September 1. Date Debit Credit
Answer #1

Question 1

Number of shares outstanding = (3000000/2) = 1500000

Book value per common share

A

Total shareholder's equity

$ 57,00,000.00

B

Number of shares outstanding

1500000

C=A/B

Book value per common share

$    3.80

Question 2

Shares to be issued as dividend (1000000 x 10%)

100000

No entry is done on the record date .

Date

Account Titles and Explanation

Debit

Credit

01-Aug

Retained earnings

$ 23,00,000.00

     Stock dividend payable

$ 23,00,000.00

(To record declaration of dividend)

Aug-20

No entry

Sep-01

Stock dividend payable

$ 23,00,000.00

      Common stock (100000 x 2)

$    2,00,000.00

       Paid in capital in excess of Par (100000 x 21)

$ 21,00,000.00

(To record distribution of stock dividend)

Stock is issued at market price and the difference between market price and Par value goes to paid in capital in excess of PAR.